OpenStudy (anonymous):

Find the intersection of 2 circles

OpenStudy (anonymous):

Circle with point A(Xa,Ya) with radius m Circle with point B(Xb,Yb) with radius n Point P(x,y) is where A and B intersect find the point(s) of intersection

OpenStudy (anonymous):

So basically, we have: (x-Xa)^2+(y-Ya)^2=m^2 (x-Xb)^2+(y-Yb)^2=n^2 and we have to solve for x and y. I kinda get lost at the plus/minus bit

OpenStudy (anonymous):

Feel free to change pronumerals to whatev suits u

OpenStudy (amistre64):

hmm, one way ... is to find the midpoint and slope between centers; define the line perp to that going thru the midpoint; then see where that hits either circle

OpenStudy (amistre64):

prolly aint the kosher solution; but its bound to get results

OpenStudy (anonymous):

Yeah, kinda looking to extending this out to find the intersection between 3 circles (triangulation)

OpenStudy (amistre64):

or ..... take waht you have, subtract the m^2 and n^2 from the sides to get them both equal to zero then equate the 2 and see what simplifies

OpenStudy (amistre64):

(x-Xa)^2+ (y-Ya)^2 - m^2 -(x-Xb)^2 - (y-Yb)^2 + n^2 -------------------------- prolly a linear in the end

OpenStudy (amistre64):

best to expand out the ^2 parts tho

OpenStudy (anonymous):

hmm... ok will try

OpenStudy (amistre64):

ewww ...... either i did something horribly wrong; or my first thought was way better lol

OpenStudy (amistre64):

is pts A and B centers of the circles? becasue now that i read it again, its a little vague

OpenStudy (anonymous):

Yeah, A and B are the centre of circles

OpenStudy (anonymous):

@amistre64, back to your first suggestion, how would finding the midpoint help?

OpenStudy (amistre64):

\[(\frac{X_a+X_b}{2},\frac{Y_a+Y_b}{2})\]

OpenStudy (amistre64):

this gives us an anchor to place the perp slope on

OpenStudy (anonymous):

But why the midpoint? Is this significant?

OpenStudy (amistre64):



OpenStudy (amistre64):

hmm, at first i thought it was ... mighta been a bad notion tho

OpenStudy (anonymous):

The midpoint won't necessarily be on the perpendicular though

OpenStudy (amistre64):

if the circles aint the same radius; i guess that defeats a midpoint

OpenStudy (amistre64):

distance between centers is the base of a triangle; with the radiuses being the other two sides ... might be applicable

OpenStudy (dumbcow):

essentially you have 2 equations with 2 variables by graphing the 2 circles you will know if you have 0,1,or 2 solutions solve this by substitution...solve each equation for y, then set them equal to get x \[\sqrt{m^{2}-(x-X_a)^{2}} + Y_a = \sqrt{n^{2}-(x-X_b)^{2}}+Y_b\] from here you can solve for x, there may be no solution lot of algebra involved